Calculus I, Fall 2021 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework week 17

1. Section 8.4, problem 48:

Figure 1: Section 8.4, problem 48.

2. Chapter 8, problem 113:

Figure 2: Chapter 8, problem 113